OpenStudy (anonymous):

Please help me find the derivative of r. r=the square root of (xsinx). Thanks!

OpenStudy (anonymous):

write it as this form (xsinx)^1/2 Try to use u-substitution and/or chain rule

OpenStudy (anonymous):

okay so this is what i have so far: 1/2(xsinx)^-1/2 * cosx

OpenStudy (anonymous):

product rule would help too: x^1/2(sinx)^1/2

OpenStudy (anonymous):

when would i apply the product rule?

OpenStudy (anonymous):

Good suggestions by @mira_14astig ... product rule here: \(\large [x^{\frac{1}{2}}\cdot (sinx)^{\frac{1}{2}}]'=[x^\frac{1}{2}]'\cdot (sinx)^\frac{1}{2}+x^\frac{1}{2}\cdot [(sinx)^\frac{1}{2}]' \)

OpenStudy (anonymous):

ohh okay i get it! thank you :)

OpenStudy (anonymous):

okay i'm still not getting the right answer...